OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Base 3 1/2feet, altitude feet. 3/4

OpenStudy (anonymous):

A = bh A = 7/2(3/4) A = 21/8

OpenStudy (ajprincess):